Session: Nov 2006 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Fri Nov 24 14:09:47 2006) Success :
No known problems. (The faulting BBC #5 in previous project has now been moved to another slot to assure it works in frequency range it 'likes'!).

Medicina (Orlati at Fri Nov 24 14:11:27 2006) Success :
everything ok.

Onsala (Michael Lindqvist at Fri Nov 24 17:42:25 2006) Success :
No known problems.

Urumqi (Zhang Hua at Fri Nov 24 18:28:48 2006) Success :

Westerbork (Tony Foley at Mon Nov 27 08:35:21 2006) Success :
Wb used in phased array mode. No known problems

Westerbork (Tony Foley at Mon Nov 27 08:55:42 2006) Minor failures : RECO
Wb used in phased array mode. Non-zero error on scan no0065 scan_check/17,el033d_ngc4151_no0065,mark4,64,2006y328d7h33m11.625s,298.6s,8.000000,16384

Effelsberg (A. Polatidis at Mon Nov 27 22:54:06 2006) Minor failures : RECO ANT POIN
Until 0911 observations were made with the subreflector at a fixed position, resulting in severe reduction in gain. Three scans of 3C303 between 0238-0358 lost due to telescope cable wrap. UT 0810-0840 lost due to work to fix the subreflector. Schedule restarted at 0923 with correct focus and pointing offsets. Calibration of the data up to 0911 will be done with corrections of the Tsys values and the new gain curve. Lost also data between UT 1321 and 1343 due to problem with the Mark 5.

Noto (P. Cassaro at Wed Nov 29 13:16:08 2006) Minor failures :
The subreflector was reported in a wrong position several times in the observation, probably with effects on the pointing.